### v2

# Autocratic strategies in Cournot oligopoly game

Abstract
An oligopoly is a market in which the price of a goods is controlled by a few firms. Cournot introduced the simplest game-theoretic model of oligopoly, where profit-maximizing behavior of each firm results in market failure. Furthermore, when the Cournot oligopoly game is infinitely repeated, firms can tacitly collude to monopolize the market. Such tacit collusion is realized by the same mechanism as direct reciprocity in the repeated prisoner’s dilemma game, where mutual cooperation can be realized whereas defection is favorable for both prisoners in one-shot game. Recently, in the repeated prisoner’s dilemma game, a class of strategies called zero-determinant strategies attracts much attention in the context of direct reciprocity. Zero-determinant strategies are autocratic strategies which unilaterally control payoffs of players. There were many attempts to find zero-determinant strategies in other games and to extend them so as to apply them to broader situations. In this paper, first, we show that zero-determinant strategies exist even in the repeated Cournot oligopoly game. Especially, we prove that an averagely unbeatable zero-determinant strategy exists, which is guaranteed to obtain the average payoff of the opponents. Second, we numerically show that the averagely unbeatable zero-determinant strategy can be used to promote collusion when it is used against an adaptively learning player, whereas it cannot promote collusion when it is used against two adaptively learning players. Our findings elucidate some negative impact of zero-determinant strategies in oligopoly market.
Author summary
Repeated games have been used to analyze the rational decision-making of multiple agents in a long-term interdependent relationship. Recently, a class of autocratic strategies, called zero-determinant strategies, was discovered in repeated games, which unilaterally controls payoffs of players. So far, properties of zero-determinant strategies in social dilemma situations have extensively been investigated, and it has been shown that some zero-determinant strategies promote cooperation. Moreover, zero-determinant strategies have been found in several games. However, it has not been known whether zero-determinant strategies exist in oligopoly games. In this paper, we investigate zero-determinant strategies in the repeated Cournot oligopoly game, which is the simplest mathematical model of oligopoly. We prove the existence of zero-determinant strategies which unilaterally enforce linear relations between the payoff of the player and the average payoff of the opponents with both positive and negative slopes. Furthermore, we numerically show that some zero-determinant strategy can promote collusion in duopoly case, although it cannot promote collusion in triopoly case. Our results imply that zero-determinant strategies can be used to promote cooperation between firms even in oligopoly market.
1 Introduction
Oligopoly is one of the simplest game-theoretic situations in economics. Cournot introduced a simple model of oligopoly, where multiple firms produce the same goods and its price decreases as the total amount of the goods increases [1, 2]. Each firm wants to maximize its profit, but needs to choose its production while considering production of the other firms. When all firms are enough rational, the Cournot-Nash equilibrium is realized, where production of each firm is the best response against production of the other firms. Because total production in this equilibrium is smaller than that in perfect competition, oligopoly results in market failure.
When the Cournot oligopoly game is infinitely repeated, situation becomes worse for consumers. Since firms become taking total profits obtained in future into account, firms can tacitly collude to monopolize the market [3], which is an example of the folk theorem in repeated games. This result is similar to direct reciprocity in the repeated prisoner’s dilemma game [4]. In the prisoner’s dilemma game, two prisoners independently choose cooperation or defection. Defection is the best action in terms of the payoff of each prisoner regardless of the action of the other prisoner, and mutual defection is realized as a result of rational behavior. Because mutual cooperation improves the payoffs of both prisoners in mutual defection, this game describes a kind of social dilemma. However, when the game is infinitely repeated, mutual cooperation can be realized as a result of long-term perspectives, similarly to tacit collusion in the repeated Cournot oligopoly game.
In 2012, a novel class of strategies, called zero-determinant (ZD) strategies, was discovered in the repeated prisoner’s dilemma game [5]. Counterintuitively, ZD strategies unilaterally control payoffs of players by enforcing linear relations between payoffs. ZD strategies can be regarded as generalization of the equalizer strategy [6], which unilaterally sets the payoff of the opponent. ZD strategies also contain the extortionate ZD strategy, by which a player never obtains payoff lower than that of the opponent, and the generous ZD strategy [7], by which a player never obtains payoff higher than that of the opponent but promotes cooperation. So far, ZD strategies were discovered in the prisoner’s dilemma game [5, 7, 8], the public goods game [9, 10], the continuous donation game [11], the asymmetric prisoner’s dilemma game [12, 13], two-player potential games [14, 15], and symmetric games with no generalized rock-paper-scissors cycles [16, 17]. Particularly, ZD strategies in games with infinite action sets are called as autocratic strategies [11]. Furthermore, ZD strategies were extended to broader situations such as games with a discount factor [18, 11, 19, 20], games with observation errors [21, 22, 23], asynchronous games [24], and stochastic games [25, 26]. Roles of ZD strategies in the context of evolution of cooperation has been extensively investigated [27, 28, 7, 29, 30, 31, 8, 32, 33].
Although the Cournot oligopoly game contains some social dilemma structure as noted above, it is not completely the same as social dilemma. This is because the action sets in the Cournot oligopoly game are infinite sets, whereas the action sets in typical social dilemma games contain only two actions, that is, cooperation and defection. Probably for that reason, ZD strategies have not been discovered in the Cournot oligopoly game, and it is not clear whether players can autocratically control payoffs in the game.
In this paper, we make two key contributions. First, we prove that ZD strategies (or, autocratic strategies) also exist in the repeated Cournot oligopoly game. Concretely, we show that ZD strategies pinning their own payoffs, positively correlated ZD strategies [34, 35], which are the class containing the extortionate ZD strategy and the generous ZD strategy, and negatively correlated ZD strategies exist in the game. Especially, we find that an averagely unbeatable ZD strategy exists in the game, by which a player unilaterally obtains the average payoff of the opponents. Second, we specify performance of the averagely unbeatable ZD strategy in evolution of collusion. We numerically investigate whether collusion is achieved when the averagely unbeatable ZD strategy is used against adaptively learning players [5, 34, 35]. We find that, in a two-player case, the averagely unbeatable ZD strategy promotes collusion of a learning player, but in a three-player case, it cannot promote collusion.
This paper is organized as follows. In Section 2, we introduce the Cournot oligopoly game and its repeated version. In Section 3, we introduce autocratic strategies, and explain their properties related to the existence. In Section 4, we prove that some types of autocratic strategies exist in the repeated Cournot oligopoly game. In Section 5, we provide numerical results. Section 6 is devoted to discussion.
2 Model
We consider the $N$-player Cournot oligopoly game with $N\geq 2$ [1, 2]. The action space of player $j$ is given by $A_{j}:=[0,x_{\mathrm{max}}]$ with $0<x_{\mathrm{max}}<\infty$ for all $j$. The action of player $j$ is described as $x_{j}\in A_{j}$, which represents production of the goods by player $j$. We collectively write the action profile as $\bm{x}:=\left(x_{1},\cdots,x_{N}\right)$. We also use the notation $\mathcal{A}:=\prod_{j}A_{j}$, $A_{-j}:=\prod_{k\neq j}A_{k}$, and $x_{-j}:=\left(x_{1},\cdots,x_{j-1},x_{j+1},\cdots,x_{N}\right)\in A_{-j}$. When we want to emphasize $x_{j}$ in $\bm{x}$, we write $\bm{x}=(x_{j},x_{-j})$. The payoff of player $j$ when the action profile is $\bm{x}$ is described as
$s_{j}\left(\bm{x}\right)$ $=\left[a-b\sum_{k=1}^{N}x_{k}\right]x_{j}\theta\left(a-b\sum_{k=1}^{N}x_{k}\right)-cx_{j},$ (1)
where $a>c>0$, $b>0$, and $\theta(\cdot)$ is the step function. That is, $\left[a-b\sum_{k=1}^{N}x_{k}\right]\theta\left(a-b\sum_{k=1}^{N}x_{k}\right)$ represents a price of the goods, and $c$ represents a cost of production. The price is a decreasing function of the total production $\sum_{k=1}^{N}x_{k}$. We assume that $x_{\mathrm{max}}\geq a/b$.
It is known that the Nash equilibrium of this game is $x_{j}=x^{(\mathrm{N})}:=\frac{a-c}{(N+1)b}$ and $s_{j}\left(\bm{x}^{(\mathrm{N})}\right)=\frac{(a-c)^{2}}{(N+1)^{2}b}$ for all $j$. On the other hand, if all players collude to monopolize the market, and share the payoffs equally, the realized state is $x_{j}=x^{(\mathrm{C})}:=\frac{a-c}{2Nb}$ and $s_{j}\left(\bm{x}^{(\mathrm{C})}\right)=\frac{(a-c)^{2}}{4Nb}$ for all $j$. Because $s_{j}\left(\bm{x}^{(\mathrm{C})}\right)>s_{j}\left(\bm{x}^{(\mathrm{N})}\right)$, the game can be regarded as one example of social dilemma. In addition, the Walras equilibrium, which corresponds to perfect competition, is $x_{j}=x^{(\mathrm{W})}:=\frac{a-c}{Nb}$ and $s_{j}\left(\bm{x}^{(\mathrm{W})}\right)=0$, where the price is equal to the marginal cost.
We consider a repeated version of the Cournot oligopoly game. We write the probability measure of an action profile at time $t$ by $P_{t}\left(\cdot\right)$. Here we consider the case that there is no discounting of future payoffs $(\delta=1)$: that is, the payoff of player $j$ in the repeated game is given by
$\mathcal{S}_{j}$ $:=\left\langle s_{j}\right\rangle^{*},$ (2)
where $\left\langle\cdots\right\rangle^{*}$ is the expected value of the quantity $\cdots$ with respect to the time-average probability of an action profile
$P^{*}\left(\cdot\right)$ $:=\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}P_{t}\left(\cdot\right).$ (3)
3 Preliminaries
The memory-one strategy of player $j$ is described as the conditional probability measure $T_{j}\left(\cdot|\bm{x}^{\prime}\right)$ of action $x_{j}$ when the action profile in the previous round is $\bm{x}^{\prime}$. We introduce the notation $s_{0}\left(\bm{x}\right):=1$. For the case of bounded payoffs $\left\{s_{k}\right\}_{k=1}^{N}$, McAvoy and Hauert introduced the concept of autocratic strategies as an extension of zero-determinant (ZD) strategies.
Definition 1 ([11]).
A memory-one strategy of player $j$ is a autocratic strategy when its strategy $T_{j}$ can be written in the form
[…]
with some coefficients $\left\{\alpha_{b}\right\}$ and some bounded function $\psi(\cdot)$.
The autocratic strategy (4) unilaterally enforces a linear relation between payoffs
00 0 $=\sum_{k=0}^{N}\alpha_{k}\left\langle s_{k}\right\rangle^{*}.$ (5)
Therefore, autocratic strategies can be used to control payoffs unilaterally. In this paper, we use two words “ZD strategy” and “autocratic strategy” interchangeably.
Below we write $B\left(\bm{x}^{\prime}\right):=\sum_{k=0}^{N}\alpha_{k}s_{k}\left(\bm{x}^{\prime}\right)$. We write the Dirac measure by $\delta_{x^{\prime}}\left(\cdot\right)$. McAvoy and Hauert proved the following proposition.
Proposition 1 ([11]).
Suppose that there exist two actions $\underline{x}_{j},\overline{x}_{j}\in A_{j}$ and $W>0$ such that
$-W$ $\leq B\left(\underline{x}_{j},x_{-j}\right)\leq 0\quad(\forall x_{-j}\in A_{-j})$
00 0 $\leq B\left(\overline{x}_{j},x_{-j}\right)\leq W\quad(\forall x_{-j}\in A_{-j}).$ (6)
Then, when we restrict the action set of player $j$ from $A_{j}$ to $A_{j}^{\prime}:=\left\{\underline{x}_{j},\overline{x}_{j}\right\}$, the memory-one strategy of player $j$
$T_{j}\left(\cdot|\bm{x}^{\prime}\right)$ $=p\left(\bm{x}^{\prime}\right)\delta_{\underline{x}_{j}}(\cdot)+\left(1-p\left(\bm{x}^{\prime}\right)\right)\delta_{\overline{x}_{j}}(\cdot)$ (7)
with
$p\left(\underline{x}_{j},x_{-j}\right)$ $:=\frac{1}{W}B\left(\underline{x}_{j},x_{-j}\right)+1\quad(\forall x_{-j}\in A_{-j})$
$p\left(\overline{x}_{j},x_{-j}\right)$ $:=\frac{1}{W}B\left(\overline{x}_{j},x_{-j}\right)\quad(\forall x_{-j}\in A_{-j})$ (8)
is an autocratic strategy unilaterally enforcing $\left\langle B\right\rangle^{*}=0$.
Indeed, we can find
[…]
with $\psi\left(\underline{x}_{j}\right)=\psi_{0}$ and $\psi\left(\overline{x}_{j}\right)=\psi_{0}+W$. McAvoy and Hauert called such autocratic strategies two-point autocratic strategies, because such autocratic strategy uses only two actions.
Although they found that the condition (6) is a sufficient condition for the existence of autocratic strategies, Ueda proved that it is also a necessary condition when the number of actions is finite [16]. However, we do not know whether it is also a necessary condition when the number of actions is infinite. At this stage, we only prove the following proposition about two-point autocratic strategies.
Proposition 2.
If a two-point autocratic strategy of player $j$ controlling $B$ exists, then there exist two actions $\underline{x}_{j},\overline{x}_{j}\in A_{j}$ such that
$B\left(\underline{x}_{j},x_{-j}\right)$ $\leq 0\quad(\forall x_{-j}\in A_{-j})$
$B\left(\overline{x}_{j},x_{-j}\right)$ $\geq 0\quad(\forall x_{-j}\in A_{-j}).$ (10)
Proof.
If a two-point autocratic strategy of player $j$ controlling $B$ exists, then its strategy satisfies
[…]
with some set $A_{j}^{\prime}=\left\{x_{j}^{(1)},x_{j}^{(2)}\right\}$ and some bounded function $\psi(\cdot)$. Because $A_{j}^{\prime}$ consists of two actions and $\psi$ is bounded, there exist
$\psi_{\mathrm{max}}:=\max_{x_{j}\in A_{j}^{\prime}}\psi(x_{j})$
$\psi_{\mathrm{min}}:=\min_{x_{j}\in A_{j}^{\prime}}\psi(x_{j})$ (12)
and
$x_{j,\mathrm{max}}:=\arg\max_{x_{j}\in A_{j}^{\prime}}\psi(x_{j})$
$x_{j,\mathrm{min}}:=\arg\min_{x_{j}\in A_{j}^{\prime}}\psi(x_{j}).$ (13)
Due to the normalization condition of probability distribution, we can rewrite Eq. (11) as
[…]
Then we find
[…]
and
[…]
because $T_{j}$ is probability. Therefore, we can regard $x_{j,\mathrm{max}}$ and $x_{j,\mathrm{min}}$ as $\underline{x}_{j}$ and $\overline{x}_{j}$, respectively. ∎
Because $B$ is bounded, the condition in Proposition 2 is equivalent to that in Proposition 1.
4 Theoretical results
In this section, we provide several examples of autocratic strategies in the Cournot oligopoly game. We remark that, in the Cournot oligopoly game, the relation
$\sum_{k\neq j}s_{k}\left(\bm{x}^{\prime}\right)$ $=\left[a-b\sum_{l=1}^{N}x_{l}\right]\sum_{k\neq j}x_{k}\theta\left(a-b\sum_{l=1}^{N}x_{l}\right)-c\sum_{k\neq j}x_{k}$ (18)
holds. Therefore, essentially, the opponents $-j$ of player $j$ can be regarded as one player, although $\sum_{k\neq j}x_{k}\in[0,(N-1)x_{\mathrm{max}}]$. Below we identify $x_{-j}$ with $\sum_{k\neq j}x_{k}$
4.1 Equalizer strategy
First, we seek for equalizer strategies of player $j$ [6, 5], that is, those unilaterally enforcing $\sum_{k\neq j}\left\langle s_{k}\right\rangle^{*}=r$ with $r\in\mathbb{R}$. For such case, we need to set $B$ as $B\left(\bm{x}^{\prime}\right):=\sum_{k\neq j}s_{k}\left(\bm{x}^{\prime}\right)-r$.
Theorem 1.
Two-point equalizer strategies in the Cournot oligopoly game do not exist for any $r$.
Proof.
According to the propositions above, two actions satisfying
$\sum_{k\neq j}s_{k}\left(\underline{x}_{j},x_{-j}\right)-r$ $\leq 0\quad(\forall x_{-j}\in A_{-j})$
$\sum_{k\neq j}s_{k}\left(\overline{x}_{j},x_{-j}\right)-r$ $\geq 0\quad(\forall x_{-j}\in A_{-j})$ (19)
are necessary for the existence of two-point equalizer strategies of player $j$. For $r>0$, when we choose $x_{-j}=0$, we find
$\sum_{k\neq j}s_{k}\left(x_{j},0\right)-r$ $=-r<0$ (20)
for any $x_{j}$. Therefore, $\overline{x}_{j}$ does not exist, and the existence condition is never satisfied. Similarly, for $r<0$, when we choose $x_{-j}=0$, we find
$\sum_{k\neq j}s_{k}\left(x_{j},0\right)-r$ $=-r>0$ (21)
for any $x_{j}$. Therefore, $\underline{x}_{j}$ does not exist, and the existence condition is never satisfied. For $r=0$, when we choose $x_{-j}=a/b$, we find
$\sum_{k\neq j}s_{k}\left(x_{j},\frac{a}{b}\right)$ $=-\frac{ca}{b}<0$ (22)
for any $x_{j}$. Therefore, $\overline{x}_{j}$ does not exist, and the existence condition is never satisfied. ∎
4.2 Strategy pinning its own payoff
Next, we seek for autocratic strategies of player $j$ pinning its own payoff, that is, those unilaterally enforcing $\left\langle s_{j}\right\rangle^{*}=r$ with $r\in\mathbb{R}$. For such case, we need to set $B$ as $B\left(\bm{x}^{\prime}\right):=s_{j}\left(\bm{x}^{\prime}\right)-r$.
Theorem 2.
Two-point autocratic strategies pinning its own payoff in the Cournot oligopoly game exist only for $-cx_{\mathrm{max}}\leq r\leq 0$.
Proof.
According to the propositions above, two actions satisfying
$s_{j}\left(\underline{x}_{j},x_{-j}\right)-r$ $\leq 0\quad(\forall x_{-j}\in A_{-j})$
$s_{j}\left(\overline{x}_{j},x_{-j}\right)-r$ $\geq 0\quad(\forall x_{-j}\in A_{-j})$ (23)
are necessary for the existence of the two-point autocratic strategies of player $j$. When $r>0$, when we choose $x_{-j}=a/b$, we find
$s_{k}\left(x_{j},\frac{a}{b}\right)-r$ $=-cx_{j}-r<0$ (24)
for any $x_{j}$. Therefore, $\overline{x}_{j}$ does not exist, and the existence condition is never satisfied.
We remark that $-cx_{\mathrm{max}}\leq s_{j}\left(\bm{x}^{\prime}\right)\leq(a-c)^{2}/(4b)$ for any $\bm{x}^{\prime}$. Therefore, the two-point autocratic strategies with $r<-cx_{\mathrm{max}}$ do not exist. When $-cx_{\mathrm{max}}\leq r\leq 0$, we find that
$s_{j}\left(0,x_{-j}\right)-r$ $=\left|r\right|\geq 0\quad(\forall x_{-j}\in A_{-j})$
$s_{j}\left(x_{\mathrm{max}},x_{-j}\right)-r$ $=-cx_{\mathrm{max}}+\left|r\right|\leq 0\quad(\forall x_{-j}\in A_{-j}).$ (25)
Therefore, according to Proposition 1, we can construct two-point autocratic strategies pinning its own payoff by using $\overline{x}_{j}=0$ and $\underline{x}_{j}=x_{\mathrm{max}}$. ∎
4.3 Positively correlated strategies
Here, we seek for positively correlated autocratic strategies of player $j$ [34, 35], which unilaterally enforce
$\left\langle s_{j}\right\rangle^{*}-\kappa$ $=\chi\left(\frac{1}{N-1}\sum_{k\neq j}\left\langle s_{k}\right\rangle^{*}-\kappa\right)$ (26)
with $\chi>0$ and $-cx_{\mathrm{max}}\leq\kappa\leq(a-c)^{2}/(4b)$. For such case, we need to set $B$ as
$B\left(\bm{x}^{\prime}\right)$ $:=s_{j}\left(\bm{x}^{\prime}\right)-\frac{\chi}{N-1}\sum_{k\neq j}s_{k}\left(\bm{x}^{\prime}\right)+(\chi-1)\kappa.$ (27)
Theorem 3.
Two-point positively correlated autocratic strategies exist only for (i) $\chi=1$ or (ii) $10<\chi<$ 1 and $\kappa\leq 0$.
Proof.
According to the propositions above, two actions satisfying
$B\left(\underline{x}_{j},x_{-j}\right)$ $\leq 0\quad(\forall x_{-j}\in A_{-j})$
$B\left(\overline{x}_{j},x_{-j}\right)$ $\geq 0\quad(\forall x_{-j}\in A_{-j})$ (28)
are necessary for the existence of two-point positively correlated autocratic strategies of player $j$.
We first rewrite the payoffs as
$s_{j}\left(\bm{x}\right)$ $=\left[\left(a-c-b\sum_{k=1}^{N}x_{k}\right)\theta\left(a-b\sum_{k=1}^{N}x_{k}\right)-c\theta\left(-\left(a-b\sum_{k=1}^{N}x_{k}\right)\right)\right]x_{j}.$ (29)
Then, $B$ is rewritten as
$B\left(\bm{x}\right)$
$=\left[\left(a-c-b\sum_{l=1}^{N}x_{l}\right)\theta\left(a-b\sum_{l=1}^{N}x_{l}\right)-c\theta\left(-\left(a-b\sum_{l=1}^{N}x_{l}\right)\right)\right]\left[x_{j}-\frac{\chi}{N-1}\sum_{k\neq j}x_{k}\right]$
$\quad+(\chi-1)\kappa.$ (30)
We find that
$B\left(\frac{\chi}{\chi+N-1}\frac{a-c}{b},x_{-j}\right)$
$=\frac{\chi}{N-1}b\left(\frac{N-1}{\chi+N-1}\frac{a-c}{b}-\sum_{k\neq j}x_{k}\right)^{2}\theta\left(c+\frac{N-1}{\chi+N-1}(a-c)-b\sum_{k\neq j}x_{k}\right)$
$\quad-c\frac{\chi}{N-1}\left(\frac{N-1}{\chi+N-1}\frac{a-c}{b}-\sum_{k\neq j}x_{k}\right)\theta\left(-\left(c+\frac{N-1}{\chi+N-1}(a-c)-b\sum_{k\neq j}x_{k}\right)\right)$
$\quad+(\chi-1)\kappa.$ (31)
The first term on the right-hand side is always non-negative. It should be noted that, when $c+(N-1)(a-c)/(\chi+N-1)-b\sum_{k\neq j}x_{k}<0$, the second term satisfies
$-c\frac{\chi}{N-1}\left(\frac{N-1}{\chi+N-1}\frac{a-c}{b}-\sum_{k\neq j}x_{k}\right)$ $>c\frac{\chi}{N-1}\frac{c}{b}>0.$ (32)
Therefore, we obtain
$B\left(\frac{\chi}{\chi+N-1}\frac{a-c}{b},x_{-j}\right)$ $\geq(\chi-1)\kappa.$ (33)
Furthermore,
$B\left(x_{\mathrm{max}},x_{-j}\right)$ $=-c\left[x_{\mathrm{max}}-\frac{\chi}{N-1}\sum_{k\neq j}x_{k}\right]+(\chi-1)\kappa.$ (34)
We also calculate
$B\left(x_{j},\frac{N-1}{\chi+N-1}\frac{a-c}{b}\right)$
$=-b\left(x_{j}-\frac{\chi}{\chi+N-1}\frac{a-c}{b}\right)^{2}\theta\left(c+\frac{\chi}{\chi+N-1}(a-c)-bx_{j}\right)$
$\quad-c\left(x_{j}-\frac{\chi}{\chi+N-1}\frac{a-c}{b}\right)\theta\left(-\left(c+\frac{\chi}{\chi+N-1}(a-c)-bx_{j}\right)\right)$
$\quad+(\chi-1)\kappa.$ (35)
The first term on the right-hand side is always non-positive. It should be noted that, when $c+\chi(a-c)/(\chi+N-1)-bx_{j}<0$, the second term satisfies
$-c\left(x_{j}-\frac{\chi}{\chi+N-1}\frac{a-c}{b}\right)$ $<-\frac{c^{2}}{b}<0.$ (36)
Therefore, we obtain
$B\left(x_{j},\frac{N-1}{\chi+N-1}\frac{a-c}{b}\right)$ $\leq(\chi-1)\kappa.$ (37)
We first consider the case $10<\chi<$ 1. When $\kappa\leq 0$, according to Eq. (33), we find
$B\left(\frac{\chi}{\chi+N-1}\frac{a-c}{b},x_{-j}\right)$ $\geq 0.$ (38)
Furthermore, by using $0\leq\sum_{k\neq j}x_{k}\leq(N-1)x_{\mathrm{max}}$ in Eq. (34), we obtain
$B\left(x_{\mathrm{max}},x_{-j}\right)$ $\leq(1-\chi)\left[\left|\kappa\right|-cx_{\mathrm{max}}\right]\leq 0,$ (39)
where we have used $-cx_{\mathrm{max}}\leq\kappa\leq(a-c)^{2}/(4b)$. Therefore, according to Proposition 1 with $\overline{x}_{j}=\chi/(\chi+N-1)\cdot(a-c)/b$ and $\underline{x}_{j}=x_{\mathrm{max}}$, two-point positively correlated autocratic strategies exist for $10<\chi<$ 1 and $\kappa\leq 0$. When $\kappa>0$, according to Eq. (37), we find
$B\left(x_{j},\frac{N-1}{\chi+N-1}\frac{a-c}{b}\right)$ $\leq(\chi-1)\kappa<0$ (40)
for any $x_{j}$. Therefore, $\overline{x}_{j}$ does not exist, and the existence condition is never satisfied.
Second, we consider the case $\chi=1$. For this case, $B$ does not depend on $\kappa$. According to Eqs. (33) and (34), we obtain
$B\left(\frac{\chi}{\chi+N-1}\frac{a-c}{b},x_{-j}\right)$ $\geq 0$ (41)
$B\left(x_{\mathrm{max}},x_{-j}\right)=-c\left[x_{\mathrm{max}}-\frac{1}{N-1}\sum_{k\neq j}x_{k}\right]$ $\leq 0.$ (42)
Therefore, according to Proposition 1 with $\overline{x}_{j}=\chi/(\chi+N-1)\cdot(a-c)/b$ and $\underline{x}_{j}=x_{\mathrm{max}}$, two-point positively correlated autocratic strategies exist for $\chi=1$.
Finally, we consider the case $\chi>1$. When $\kappa\geq 0$, from Eq. (30), we find
$B\left(x_{j},(N-1)x_{\mathrm{max}}\right)$ $=-c\left[x_{j}-\chi x_{\mathrm{max}}\right]+(\chi-1)\kappa>0$ (43)
for any $x_{j}$. Therefore, $\underline{x}_{j}$ does not exist, and the existence condition is never satisfied. When $\kappa<0$, according to Eq. (37), we obtain
$B\left(x_{j},\frac{N-1}{\chi+N-1}\frac{a-c}{b}\right)$ $\leq(\chi-1)\kappa<0$ (44)
for any $x_{j}$. Therefore, $\overline{x}_{j}$ does not exist, and the existence condition is never satisfied. ∎
Because $\chi\leq 1$, these positively correlated autocratic strategies may be interpreted as generous strategies [7] rather than extortionate strategies [5]. We remark that, when $\chi\rightarrow 0$, the two-point positively correlated autocratic strategy is reduced to the autocratic strategy in Theorem 2, since $\overline{x}_{j}=\chi/(\chi+N-1)\cdot(a-c)/b\rightarrow 0$. Moreover, when $\chi=1$, the two-point positively correlated autocratic strategy can be regarded as an averagely unbeatable strategy, since it unilaterally enforces
$\left\langle s_{j}\right\rangle^{*}$ $=\frac{1}{N-1}\sum_{k\neq j}\left\langle s_{k}\right\rangle^{*}.$ (45)
When $N=2$, it is reduced to an unbeatable strategy [36, 37, 14, 16]. Note that, when $\chi=1$, the action $\overline{x}_{j}=\chi/(\chi+N-1)\cdot(a-c)/b=(a-c)/(Nb)$ in Eq. (41) is one in the Walras equilibrium $x^{(\mathrm{W})}$. It has been known that the Walras equilibrium action $x^{(\mathrm{W})}$ is an unbeatable action against any monomorphic opponents [38]. The averagely unbeatable autocratic strategy uses this property of the Walras equilibrium action to control payoffs in the repeated game.
4.4 Negatively correlated strategies
Here, we seek for negatively correlated autocratic strategies of player $j$, which unilaterally enforce
$\left\langle s_{j}\right\rangle^{*}-\kappa$ $=\chi\left(\frac{1}{N-1}\sum_{k\neq j}\left\langle s_{k}\right\rangle^{*}-\kappa\right)$ (46)
with $\chi<0$ and $-cx_{\mathrm{max}}\leq\kappa\leq(a-c)^{2}/(4b)$. For such case, we again need to set $B$ as
$B\left(\bm{x}^{\prime}\right)$ $:=s_{j}\left(\bm{x}^{\prime}\right)-\frac{\chi}{N-1}\sum_{k\neq j}s_{k}\left(\bm{x}^{\prime}\right)+(\chi-1)\kappa.$ (47)
Theorem 4.
Two-point negatively correlated autocratic strategies exist only for $\left|\chi\right|\leq 1$ and $-cx_{\mathrm{max}}/\left(1+\left|\chi\right|\right)\leq\kappa\leq-\left|\chi\right|cx_{\mathrm{max}}/\left(1+\left|\chi\right|\right)$.
Proof.
According to the propositions above, two actions satisfying
$B\left(\underline{x}_{j},x_{-j}\right)$ $\leq 0\quad(\forall x_{-j}\in A_{-j})$
$B\left(\overline{x}_{j},x_{-j}\right)$ $\geq 0\quad(\forall x_{-j}\in A_{-j})$ (48)
are necessary for the existence of two-point negatively correlated autocratic strategies of player $j$. We remark that $B$ can be explicitly written as
$B\left(\bm{x}\right)$ $=\left[\left(a-b\sum_{l=1}^{N}x_{l}\right)\theta\left(a-b\sum_{l=1}^{N}x_{l}\right)-c\right]\left[x_{j}+\frac{\left|\chi\right|}{N-1}\sum_{k\neq j}x_{k}\right]-\left(\left|\chi\right|+1\right)\kappa.$ (49)
When $\kappa>-\left|\chi\right|cx_{\mathrm{max}}/\left(1+\left|\chi\right|\right)$ with any $\chi<0$, we find
$B\left(x_{j},(N-1)x_{\mathrm{max}}\right)$ $=-c\left[x_{j}+\left|\chi\right|x_{\mathrm{max}}\right]-\left(\left|\chi\right|+1\right)\kappa$
$<-c\left[x_{j}+\left|\chi\right|x_{\mathrm{max}}\right]+\left|\chi\right|cx_{\mathrm{max}}$
$=-cx_{j}$
$\leq 0$ (50)
for all $x_{j}$. Therefore, $\overline{x}_{j}$ does not exist.
When $\left|\chi\right|>1$ and $-cx_{\mathrm{max}}\leq\kappa\leq-\left|\chi\right|cx_{\mathrm{max}}/\left(1+\left|\chi\right|\right)$, we find
$B\left(x_{j},0\right)$ $=\left[(a-bx_{j})\theta(a-bx_{j})-c\right]x_{j}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$\geq-cx_{\mathrm{max}}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$\geq-cx_{\mathrm{max}}+\left|\chi\right|cx_{\mathrm{max}}$
$>0$ (51)
for all $x_{j}$. Therefore, $\underline{x}_{j}$ does not exist.
When $\left|\chi\right|\leq 1$ and $-cx_{\mathrm{max}}/\left(1+\left|\chi\right|\right)\leq\kappa\leq-\left|\chi\right|cx_{\mathrm{max}}/\left(1+\left|\chi\right|\right)$, we find
$B\left(0,x_{-j}\right)$ $=\left[\left(a-b\sum_{k\neq j}x_{k}\right)\theta\left(a-b\sum_{k\neq j}x_{k}\right)-c\right]\frac{\left|\chi\right|}{N-1}\sum_{k\neq j}x_{k}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$\geq-\left|\chi\right|cx_{\mathrm{max}}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$\geq 0.$ (52)
Therefore, we can use $x_{j}=0$ as $\overline{x}_{j}$. Furthermore, we also find
$B\left(x_{\mathrm{max}},x_{-j}\right)$ $=-cx_{\mathrm{max}}-c\frac{\left|\chi\right|}{N-1}\sum_{k\neq j}x_{k}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$\leq-cx_{\mathrm{max}}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$\leq 0.$ (53)
Therefore, we can use $x_{j}=x_{\mathrm{max}}$ as $\underline{x}_{j}$. According to Proposition 1, two-point negatively correlated autocratic strategies exist for this parameter region.
Finally, when $\left|\chi\right|\leq 1$ and $-cx_{\mathrm{max}}\leq\kappa<-cx_{\mathrm{max}}/\left(1+\left|\chi\right|\right)$, we find
$B\left(x_{j},0\right)$ $=\left[(a-bx_{j})\theta(a-bx_{j})-c\right]x_{j}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$\geq-cx_{\mathrm{max}}+\left(\left|\chi\right|+1\right)\left|\kappa\right|$
$>-cx_{\mathrm{max}}+cx_{\mathrm{max}}$
$=0$ (54)
for all $x_{j}$. Therefore, $\underline{x}_{j}$ does not exist. ∎
It should be noted that when $\chi\rightarrow-0$, the two-point negatively correlated autocratic strategy is also reduced to the autocratic strategy in Theorem 2. When $\chi=-1$, the two-point negatively correlated autocratic strategy is similar to the ZD strategy pinning the sum of payoffs of two players in the prisoner’s dilemma game [18, 39].
5 Numerical results
In this section, we numerically investigate performance of the autocratic strategies in the previous section. We set model parameters $a=2.0$, $b=1.0$, $c=1.0$, and $x_{\mathrm{max}}=2.5$. Because $B$ in Eqs. (27) and (47) for the two-point autocratic strategies satisfy
$\left|B\left(\bm{x}^{\prime}\right)\right|$ $\leq\left|s_{j}\left(\bm{x}^{\prime}\right)\right|+\frac{\left|\chi\right|}{N-1}\sum_{k\neq j}\left|s_{k}\left(\bm{x}^{\prime}\right)\right|+\left|(\chi-1)\kappa\right|$
$\leq(1+\left|\chi\right|)\max\left\{\frac{(a-c)^{2}}{4b},cx_{\mathrm{max}}\right\}+\left|(\chi-1)\kappa\right|,$ (55)
we set
$W$ $=(1+\left|\chi\right|)\max\left\{\frac{(a-c)^{2}}{4b},cx_{\mathrm{max}}\right\}+\left|(\chi-1)\kappa\right|.$ (56)
We approximately calculate the payoffs $\mathcal{S}_{k}$ in the repeated game by using time average of one sample: $\sum_{t=1}^{N_{\mathrm{TIME}}}s_{k}^{(t)}/N_{\mathrm{TIME}}$.
5.1 Autocratic strategies against fixed memory-zero opponents
First, we consider situations where player $j$ uses the two-point autocratic strategies with various $(\chi,\kappa)$ and all opponents repeat fixed actions. We set $N=10$ and $N_{\mathrm{TIME}}=10^{6}$. We assume that all other players $k\neq j$ use the same memory-zero strategy
$T_{k}\left(\cdot|\bm{x}^{\prime}\right)$ $=\delta_{x_{k}}(\cdot)$ (57)
with $x_{k}=0.01\times i_{\mathrm{oppo}}/(N-1)$, and change the integer $i_{\mathrm{oppo}}$ in the range $i_{\mathrm{oppo}}\in[0,200]$. The procedure is summarized in Algorithm 1.
Algorithm 1 Autocratic strategies against fixed memory-zero opponents
0:  Parameters of models $a,b,c,x_{\mathrm{max}},N$, total time $N_{\mathrm{TIME}}$
0:  Feasible parameters of two-point autocratic strategies $\chi,\kappa,W$
0:  Time-averaged payoffs $S_{l}$ for all $l$
1:  Set $\overline{x}_{j}$ and $\underline{x}_{j}$
2:  for $i_{\mathrm{oppo}}=0$ to $200$ do
3:     Initial condition: $x_{j}=\overline{x}_{j}$
4:     for $k\neq j$ do
5:        $x_{k}=0.01\times i_{\mathrm{oppo}}/(N-1)$
6:     end for
7:     Initialize time-averages payoffs $S_{l}\leftarrow 0$ for all $l$
8:     Calculate $B$ from Eq. (27) or (47)
9:     for $t=1$ to $N_{\mathrm{TIME}}$ do
10:        Update $x_{j}$ by transition probability (7) and (8)
11:        Calculate payoffs (1) for all players
12:        Update total payoffs $S_{l}\leftarrow S_{l}+s_{l}$ for all $l$
13:        Calculate $B$ from Eq. (27) or (47)
14:     end for
15:     Output time-averaged payoffs $S_{l}/N_{\mathrm{TIME}}$ for all $l$
16:  end for
In Fig. 1, we display a relation between $\mathcal{S}_{j}$ and $\sum_{k\neq j}\mathcal{S}_{k}/(N-1)$ for various $(\chi,\kappa)$.
Fig 1: A linear relation between $\mathcal{S}_{j}$ and $\sum_{k\neq j}\mathcal{S}_{k}/(N-1)$ when player $j$ uses the two-point autocratic strategies with various $(\chi,\kappa)$ and all opponents repeat fixed actions.
We find that a linear relation (26) or (46) is indeed enforced.
5.2 Autocratic strategies against random memory-one opponents
Next, we consider the situation where player $j$ uses a two-point autocratic strategy with $(\chi,\kappa)=(1,0)$, and the opponents $k\neq j$ adopt random memory-one strategies. We assume that the opponents only take the cooperative action $x^{(\mathrm{C})}$ or the Nash equilibrium action $x^{(\mathrm{N})}$. A memory-one strategy of player $k$ is written as
$T_{k}\left(\cdot|\bm{x}^{\prime}\right)$ $=q_{k}\left(\bm{x}^{\prime}\right)\delta_{x^{(\mathrm{C})}}(\cdot)+\left(1-q_{k}\left(\bm{x}^{\prime}\right)\right)\delta_{x^{(\mathrm{N})}}(\cdot),$ (58)
where $q_{k}\left(\bm{x}^{\prime}\right)$ corresponds to the probability to use $x^{(\mathrm{C})}$ when the action profile in the previous round was $\bm{x}^{\prime}$. For simplicity, we below write $\underline{x}_{j}$ and $\overline{x}_{j}$ as $\underline{x}$ and $\overline{x}$, respectively. We represent the memory-one strategies by a vector [9]
$\bm{q}_{k}$ $=\left(q_{\underline{x},x^{(\mathrm{C})},0},\dots,q_{\underline{x},x^{(\mathrm{C})},N-2},\ q_{\underline{x},x^{(\mathrm{N})},0},\dots,q_{\underline{x},x^{(\mathrm{N})},N-2},\right.$
$\qquad\left.q_{\overline{x},x^{(\mathrm{C})},0},\dots,q_{\overline{x},x^{(\mathrm{C})},N-2},\ q_{\overline{x},x^{(\mathrm{N})},0},\dots,q_{\overline{x},x^{(\mathrm{N})},N-2}\right).$ (59)
The entries $q_{x_{j},x_{k},i}$ denote the probability to take $x^{(\mathrm{C})}$ in the next round, given that the autocratic strategy previously played $x_{j}\in\{\underline{x},\overline{x}\}$, player $k$ previously played $x_{k}\in\{x^{(\mathrm{C})},x^{(\mathrm{N})}\}$, and that $i$ of the other $N-2$ players took $x^{(\mathrm{C})}$. We randomly generate $200$ memory-one strategy profiles $\{\bm{q}_{k}\}_{k\neq j}$ from a uniform distribution. The procedure is summarized in Algorithm 2. We again set $N=10$ and $N_{\mathrm{TIME}}=10^{6}$.
Algorithm 2 Autocratic strategies against random memory-one opponents
0:  Parameters of models $a,b,c,x_{\mathrm{max}},N$, total time $N_{\mathrm{TIME}}$
0:  Feasible parameters of two-point autocratic strategies $\chi,\kappa,W$
0:  Time-averaged payoffs $S_{l}$ for all $l$
1:  Set $\overline{x}_{j}$ and $\underline{x}_{j}$
2:  Set $x^{(\mathrm{C})}$ and $x^{(\mathrm{N})}$
3:  for $i_{\mathrm{oppo}}=1$ to $200$ do
4:     Initial condition: $x_{j}=\overline{x}_{j}$
5:     for $k\neq j$ do
6:        Initialize each normal agent’s strategy vector $\bm{q}_{k}\in[0,1]^{4(N-1)}$ $(k\neq j)$ randomly
7:        Initial condition: $x_{k}=x^{(\mathrm{C})}$
8:     end for
9:     Initialize time-averages payoffs $S_{l}\leftarrow 0$ for all $l$
10:     Calculate $B$ from Eq. (27) or (47)
11:     for $t=1$ to $N_{\mathrm{TIME}}$ do
12:        Update $x_{j}$ by transition probability (7) and (8)
13:        for $k\neq j$ do
14:           Update $x_{k}$ by transition probability (59)
15:        end for
16:        Calculate payoffs (1) for all players
17:        Update total payoffs $S_{l}\leftarrow S_{l}+s_{l}$ for all $l$
18:        Calculate $B$ from Eq. (27) or (47)
19:     end for
20:     Output time-averaged payoffs $S_{l}/N_{\mathrm{TIME}}$ for all $l$
21:  end for
Figure 2 shows a relationship between the payoff of the autocratic player (horizontal) and the average payoffs of $N-1$ players (vertical). We find that the autocratic strategy unilaterally enforces a linear relationship between payoffs. As shown in Fig. 1 and Fig. 2, this holds regardless of memory length (0 or 1) or how the opponents’ strategies are defined, whether using fixed values with a certain separation or randomized memory-one rules. The two-point autocratic strategy consistently enforces the intended linear relation under all these conditions.
Fig 2: A linear relation between $\mathcal{S}_{j}$ and $\sum_{k\neq j}\mathcal{S}_{k}/(N-1)$ when player $j$ uses a two-point autocratic strategy with $(\chi,\kappa)=(1,0)$ against random memory-one strategies.
5.3 Autocratic strategies against adaptive memory-one opponents
Finally, we investigate the scenario in which memory-one opponents attempt to maximize their payoffs against an autocratic strategy. We refer to such players as adaptive learning players. In the repeated prisoner’s dilemma game, it has been shown that the positively correlated ZD strategy can force an adaptive learning player to cooperate unconditionally [5, 34, 35]. Here we again assume that memory-one strategies of the opponents take the form (59).
To numerically implement this setting, we first initialize actions of all agents and the opponents’ memory-one strategies with random values. The opponents then adaptively update their memory-one strategies using a greedy method. Specifically, at each round, an opponent selects one of its strategy parameters $q$ (corresponding to a given previous action profile), and perturbs it by either $+0.001$ or $-0.001$. Both the original $q$ and the perturbed $\tilde{q}$ are used to determine the action $x_{k}$ and $\tilde{x}_{k}$, respectively. Then, the payoff is calculated against the other players. If the perturbation leads to a higher payoff, the change is accepted; otherwise, it is rejected. This process is repeated iteratively, allowing each opponent to gradually improve its strategy against the other players. The detailed algorithm is provided in Algorithm 3. We set $N=2$ for Fig. 3 and $N=3$ for Fig. 5. Although $N_{\mathrm{TIME}}=10^{8}$ was used, plotting was terminated once no further payoff improvements were observed.
Algorithm 3 Autocratic strategies against adaptive memory-one opponents
0:  Parameters of models $a,b,c,x_{\mathrm{max}},N$, total time $N_{\mathrm{TIME}}$
0:  Feasible parameters of two-point autocratic strategies $\chi,\kappa,W$
0:  Time-averaged payoffs $S_{l}$ for all $l$ and average strategies $\sum_{k\neq j}\bm{q}_{k}/(N-1)$
1:  Set $\overline{x}_{j}$ and $\underline{x}_{j}$
2:  Set $x^{(\mathrm{C})}$ and $x^{(\mathrm{N})}$
3:  Initial condition: $x_{j}\leftarrow$ randomly choose $\overline{x}$ or $\underline{x}$
4:  for $k\neq j$ do
5:     Initialize each normal agent’s strategy vector $\bm{q}_{k}\in[0,1]^{4(N-1)}$ $(k\neq j)$ randomly
6:     Initial condition: $x_{k}\leftarrow$ randomly choose $x^{(\mathrm{C})}$ or $x^{(\mathrm{N})}$
7:  end for
8:  Initialize time-averages payoffs $S_{l}\leftarrow 0$ for all $l$
9:  Calculate $B$ from Eq. (27) or (47)
10:  for $t=1$ to $N_{\text{TIME}}$ do
11:     for $k\neq j$ do
12:        Sample action $x_{k}$ with probability $q_{x_{j}^{\prime},x_{k}^{\prime},i^{\prime}}$ corresponding to previous action profile $\bm{x}^{\prime}$
13:        Create perturbed strategy $\tilde{q}_{x_{j}^{\prime},x_{k}^{\prime},i^{\prime}}\leftarrow\text{clip}(q_{x_{j}^{\prime},x_{k}^{\prime},i^{\prime}}+\delta,0,1)$ with $\delta=\pm 0.001$
14:        Sample alternative action $\tilde{x}_{k}$ using $\tilde{q}_{x_{j}^{\prime},x_{k}^{\prime},i^{\prime}}$
15:        Calculate payoffs $s_{k}$ (from $x_{k}$) and $\tilde{s}_{k}$ (from $\tilde{x}_{k}$)
16:        if $\tilde{s}_{k}>s_{k}$ then
17:           Update $q_{x_{j}^{\prime},x_{k}^{\prime},i^{\prime}}\leftarrow\tilde{q}_{x_{j}^{\prime},x_{k}^{\prime},i^{\prime}}$, adopt $x_{k}\leftarrow\tilde{x}_{k}$
18:        else
19:           Keep $q_{x_{j}^{\prime},x_{k}^{\prime},i^{\prime}}$ and $x_{k}$
20:        end if
21:     end for
22:     Update $x_{j}$ by transition probability (7) and (8)
23:     Calculate payoffs (1) for all players
24:     Update total payoffs $S_{l}\leftarrow S_{l}+s_{l}$ for all $l$
25:     Calculate $B$ from Eq. (27) or (47)
26:     Output time-averaged payoffs $S_{l}/t$ for all $l$ and average strategies $\sum_{k\neq j}\bm{q}_{k}/(N-1)$
27:  end for
Figure 3 shows the time evolution of the strategy used by the adaptive memory-one player against a fixed autocratic strategy in the case of $N=2$, while Fig. 4 shows the time evolution of the time-averaged payoffs of both players.
Due to the linear payoff relationship between the two players and the fact that this relationship is unilaterally enforced by the autocratic strategy, an adaptive memory-one player attempting to maximize its own payoff is inevitably driven to adopt the cooperative action $x^{(\mathrm{C})}$, as shown in [5, 34, 35].
In fact, as time progresses, the adaptive player’s time-averaged payoff increases and gradually approaches the theoretical expected payoff of $0.05405$ obtained when it consistently chooses $x^{(\mathrm{C})}$ (Fig. 4; see S1 Appendix. SI1 for derivation).
Once the payoff converges to this value, we observe that $q_{\underline{x},x^{(\mathrm{C})},0}=1$ and $q_{\overline{x},x^{(\mathrm{C})},0}=1$, indicating that the adaptive memory-one player chooses $x^{(\mathrm{C})}$ with probability one in all subsequent rounds.
Fig 3: Strategy adaptation of an adaptive memory-one player against an autocratic strategy in the duopoly case ($N=2$).
Fig 4: Time-averaged payoffs of a fixed autocratic player and an adaptive memory-one player in the duopoly case ($N=2$). Data points are plotted every 1000 time steps.
Figure 5 shows the time evolution of the strategy used by the adaptive memory-one players against a fixed autocratic strategy in the case of $N=3$, while Fig. 6 shows the time evolution of the time-averaged average payoffs of both players.
According to Fig. 6, in the case of $N=3$, the payoffs of the two adaptive memory-one players do not continue to increase but rather converge around 0.04015. Notably, this value is significantly lower than the theoretical expected payoff of 0.05039, which would be achieved if both adaptive players consistently chose $x^{(\mathrm{C})}$ (see S1 Appendix. SI1).
This result indicates that, in the case of $N=3$, the adaptive players do not evolve their strategies toward cooperation. Indeed, as shown in Fig. 6, the strategy converges to $\sum q_{\overline{x},x^{(\mathrm{C})},1}/2=0$, $\sum q_{\overline{x},x^{(\mathrm{N})},0}/2=0$, $\sum q_{\underline{x},x^{(\mathrm{C})},1}/2=1$, and $\sum q_{\underline{x},x^{(\mathrm{N})},0}/2=1$, suggesting that the two adaptive players evolve their strategies in a synchronized manner.
In other words, rather than attempting to maximize their payoffs against the fixed autocratic strategy, the adaptive random memory-one players appear to have evolved strategies that optimize their payoffs against each other. A more detailed discussion of this phenomenon is provided in Subsection 6.2.
Fig 5: Strategy adaptation of two adaptive memory-one players against an autocratic strategy in the triopoly case ($N=3$).
Fig 6: Time-averaged payoffs of a fixed autocratic player and two adaptive memory-one players in the triopoly case ($N=3$). Data points are plotted every 1000 time steps.
6 Discussion
6.1 Difference between Cournot oligopoly and prisoner’s dilemma
In this paper, we found several autocratic strategies in the repeated Cournot oligopoly game. As noted in Section 2, when action sets of all players are restricted to $\{x^{(\mathrm{C})},x^{(\mathrm{N})}\}$ and $N=2$, the Cournot oligopoly game is reduced to the prisoner’s dilemma game, where $x^{(\mathrm{C})}$ and $x^{(\mathrm{N})}$ correspond cooperation and defection, respectively (see S1 Appendix. SI2). However, when action sets are $[0,x_{\mathrm{max}}]$, resulting autocratic strategies are quite different from those in the the prisoner’s dilemma game. In the the prisoner’s dilemma case, the equalizer strategies exist [6, 5], contrary to our Theorem 1. Similarly, although the autocratic strategies pinning its own payoff do not exist in the prisoner’s dilemma game [5], they exist in the Cournot oligopoly game (Theorem 2). Furthermore, whereas our Theorem 3 claims that two-point positively correlated autocratic strategies exist only for $\chi\leq 1$, positively correlated autocratic strategies exist for $\chi\geq 1$ for the prisoner’s dilemma game [5, 7]. These differences also exist even if we compare the Cournot oligopoly game with the public goods game [9, 10], which is an $N$-player version of the prisoner’s dilemma game. These differences seem to come from the nature of the payoff function of the Cournot oligopoly game.
Obviously, these differences in equalizer strategies and positively correlated autocratic strategies may come from the fact that we only consider two-point autocratic strategies. However, it is known that, when action sets of all players are finite sets, the existence condition of two-point zero-determinant strategies is equivalent to the existence condition of general zero-determinant strategies [16, 15]. We expect that this equivalence also holds when an action set of some player is an infinite set. Specifying a necessary and sufficient condition for the existence of zero-determinant strategies in the case that action sets of some players are infinite sets is an important future problem.
6.2 Interpretation of numerical results
In Section 5, we found that the averagely unbeatable autocratic strategy can promotes collusion of adaptively learning players in duopoly ($N=2$) case, but cannot promote collusion in triopoly ($N=3$) case. The result in duopoly case is similar to that in the repeated prisoner’s dilemma game [5, 34, 35]. In the prisoner’s dilemma game, the extortionate ZD strategy can force an adaptively learning player (without a theory of mind) to cooperate unconditionally. This comes from a property of positively correlated ZD strategies: increase in payoff of the opponent implies increase in payoff of a player using the ZD strategies. However, in triopoly case, the averagely unbeatable autocratic strategy cannot force the adaptive opponents to collude. This means that enforcing a linear relation (45) is not sufficient to directly control all other opponents; the averagely unbeatable autocratic strategy can control only the average payoff of the opponents.
The failure of control by the averagely unbeatable autocratic strategy in triopoly case may come from the nature of two adaptive agents. We provide the values of one-shot payoffs for $N=2$ and $N=3$ in Table 1 and Table 2, respectively. When two adaptive agents regard an autocratic agent as an environment [40], they selfishly learn their best response $x^{(\mathrm{N})}$, since $x^{(\mathrm{N})}$ brings an agent larger one-shot payoff than $x^{(\mathrm{C})}$ as long as an autocratic agent takes $\overline{x}_{j}$. (It should be noted that an autocratic agent frequently takes $\overline{x}_{j}$ and rarely takes $\underline{x}_{j}$ in our parameters.) Concretely, in Table 2, when player $1$ is an autocratic agent, the payoff of player $2$ satisfies $s_{2}\left(\overline{x},x^{(\mathrm{C})},x^{(\mathrm{C})}\right)<s_{2}\left(\overline{x},x^{(\mathrm{N})},x^{(\mathrm{C})}\right)$ and $s_{2}\left(\overline{x},x^{(\mathrm{C})},x^{(\mathrm{N})}\right)=s_{2}\left(\overline{x},x^{(\mathrm{N})},x^{(\mathrm{N})}\right)$. Therefore, they come to repeat $x^{(\mathrm{N})}$ for most rounds. Only when the state of “environment” $x_{j}$ switches to $\underline{x}_{j}$, they synchronously take $x^{(\mathrm{C})}$, since $s_{2}\left(\underline{x},x^{(\mathrm{C})},x^{(\mathrm{C})}\right)>s_{2}\left(\underline{x},x^{(\mathrm{N})},x^{(\mathrm{C})}\right)$ and $s_{2}\left(\underline{x},x^{(\mathrm{C})},x^{(\mathrm{N})}\right)>s_{2}\left(\underline{x},x^{(\mathrm{N})},x^{(\mathrm{N})}\right)$ in Table 2. This is as if they synchronously take $x^{(\mathrm{C})}$ in order to bring back the state of “environment” from $\underline{x}_{j}$ to $\overline{x}_{j}$, because transition between states of “environment” occurs with probability proportional to $\left|B(\bm{x}^{\prime})\right|$. In S1 Appendix. SI2, we provide a numerical result in a situation where only two adaptive agents exist. Both adaptive agents finally learned to repeat the Nash equilibrium action $x^{(\mathrm{N})}$. The failure of control in triopoly case seems to come from the same reason as this situation. It should be noted that a similar phenomenon occurs in the repeated public goods game [41], where a cooperation enforcing strategy is adopted against two independent agents using reinforcement learning. The cooperation enforcing strategy is known to be an unbeatable ZD strategy [17]. Robustness of our numerical result against other learning algorithms should be investigated in future.
{adjustwidth}
-2.25in0in     $x^{(\mathrm{C})}$ $x^{(\mathrm{N})}$ $\overline{x}$     $\frac{1}{8}\frac{(a-c)^{2}}{b},\frac{1}{16}\frac{(a-c)^{2}}{b}$ $\frac{1}{12}\frac{(a-c)^{2}}{b},\frac{1}{18}\frac{(a-c)^{2}}{b}$ $\underline{x}$     $-cx_{\mathrm{max}},-\frac{1}{4}\frac{c(a-c)}{b}$ $-cx_{\mathrm{max}},-\frac{1}{3}\frac{c(a-c)}{b}$
Table 1: Payoffs for $N=2$. Player $1$ is an autocratic agent, and player $2$ takes either $x^{(\mathrm{C})}$ or $x^{(\mathrm{N})}$.
{adjustwidth}
-2.25in0in     $\left(x^{(\mathrm{C})},x^{(\mathrm{C})}\right)$ $\left(x^{(\mathrm{C})},x^{(\mathrm{N})}\right)$ $\left(x^{(\mathrm{N})},x^{(\mathrm{C})}\right)$ $\left(x^{(\mathrm{N})},x^{(\mathrm{N})}\right)$ $\overline{x}$     $\frac{1}{9}\frac{(a-c)^{2}}{b},\frac{1}{18}\frac{(a-c)^{2}}{b},\frac{1}{18}\frac{(a-c)^{2}}{b}$ $\frac{1}{12}\frac{(a-c)^{2}}{b},\frac{1}{24}\frac{(a-c)^{2}}{b},\frac{1}{16}\frac{(a-c)^{2}}{b}$ $\frac{1}{12}\frac{(a-c)^{2}}{b},\frac{1}{16}\frac{(a-c)^{2}}{b},\frac{1}{24}\frac{(a-c)^{2}}{b}$ $\frac{1}{18}\frac{(a-c)^{2}}{b},\frac{1}{24}\frac{(a-c)^{2}}{b},\frac{1}{24}\frac{(a-c)^{2}}{b}$ $\underline{x}$     $-cx_{\mathrm{max}},-\frac{1}{6}\frac{c(a-c)}{b},-\frac{1}{6}\frac{c(a-c)}{b}$ $-cx_{\mathrm{max}},-\frac{1}{6}\frac{c(a-c)}{b},-\frac{1}{4}\frac{c(a-c)}{b}$ $-cx_{\mathrm{max}},-\frac{1}{4}\frac{c(a-c)}{b},-\frac{1}{6}\frac{c(a-c)}{b}$ $-cx_{\mathrm{max}},-\frac{1}{4}\frac{c(a-c)}{b},-\frac{1}{4}\frac{c(a-c)}{b}$
Table 2: Payoffs for $N=3$. Player $1$ is an autocratic agent, and players $2$ and $3$ take either $x^{(\mathrm{C})}$ or $x^{(\mathrm{N})}$.
Although the averagely unbeatable autocratic strategy cannot promote collusion of adaptive agents for $N=3$ (and probably for $N\geq 4$), our result for $N=2$ suggests another way to collusion. As a special property of the Cournot oligopoly game, players $-j$ can be regarded as one effective player, as shown in Eq. (18). Therefore, when players $-j$ use a ZD alliance [9], which is a ZD strategy implemented by a group of players, they can lead the opponent $j$ to collusion. In S1 Appendix. SI3, we prove the existence of an averagely unbeatable ZD alliance in the Cournot oligopoly game. We expect that the averagely unbeatable ZD alliance by $N-1$ players can promote collusion of the other one player.
Supporting information
S1 Appendix.
SI1 Calculation of expected payoffs when the opponents use $x^{(\mathrm{C})}$
In this appendix, we calculate the expected payoffs of players when player $j$ uses the positively-correlated autocratic strategy with $\chi=1$ in Section 5 and all other players $k\neq j$ use a constant action $x_{k}=x^{(\mathrm{C})}$. For such case, the time evolution is effectively a two-state Markov chain of $x_{j}$. That is, the time evolution of the probability distribution of $x_{j}$ at $t$-th round, $P_{t}(x_{j})$, is described as
$\left(\begin{array}[]{c}P_{t+1}\left(\overline{x}_{j}\right)\\ P_{t+1}\left(\underline{x}_{j}\right)\end{array}\right)$ $=\left(\begin{array}[]{cc}1-\mu&\nu\\ \mu&1-\nu\end{array}\right)\left(\begin{array}[]{c}P_{t}\left(\overline{x}_{j}\right)\\ P_{t}\left(\underline{x}_{j}\right)\end{array}\right)$ (SI7)
with $\overline{x}_{j}=(a-c)/(Nb)$ and $\underline{x}_{j}=x_{\mathrm{max}}$, where we have introduced
$\mu$ $:=p\left(\overline{x}_{j},(N-1)x^{(\mathrm{C})}\right)=\frac{1}{W}\frac{(N-1)(a-c)^{2}}{4N^{2}b}$ (SI8)
$\nu$ $:=1-p\left(\underline{x}_{j},(N-1)x^{(\mathrm{C})}\right)=\frac{1}{W}c\left(x_{\mathrm{max}}-\frac{a-c}{2Nb}\right)$ (SI9)
according to Eq. (7). The stationary probability distribution of the Markov chain is calculated as
$P^{(\mathrm{s})}\left(\overline{x}_{j}\right)$ $=\frac{\nu}{\mu+\nu}$
$P^{(\mathrm{s})}\left(\underline{x}_{j}\right)$ $=\frac{\mu}{\mu+\nu}.$ (SI10)
Then, the expected payoff of player $j$ is obtained as
$\left\langle s_{j}\right\rangle^{*}$ $=P^{(\mathrm{s})}\left(\overline{x}_{j}\right)s_{j}\left(\overline{x}_{j},(N-1)x^{(\mathrm{C})}\right)+P^{(\mathrm{s})}\left(\underline{x}_{j}\right)s_{j}\left(\underline{x}_{j},(N-1)x^{(\mathrm{C})}\right)$
$=\frac{\frac{(N-1)(a-c)^{2}}{4N^{2}b}c\left(x_{\mathrm{max}}-\frac{a-c}{Nb}\right)}{\frac{(N-1)(a-c)^{2}}{4N^{2}b}+c\left(x_{\mathrm{max}}-\frac{a-c}{2Nb}\right)}.$ (SI11)
The expected payoffs of other players are equal to $\left\langle s_{k}\right\rangle^{*}=\left\langle s_{j}\right\rangle^{*}$ for all $k\neq j$, because the autocratic strategy of player $j$ unilaterally enforces Eq. (45) and all other players use the same constant action $x^{(\mathrm{C})}$. For the values of parameters used in Section 5, we estimate $\left\langle s_{j}\right\rangle^{*}\simeq 0.05405$ for $N=2$, and $\left\langle s_{j}\right\rangle^{*}\simeq 0.05039$ for $N=3$.
SI2 Numerical results for only two adaptive agents
In this appendix, we present additional numerical results involving only two adaptive memory-one players, in order to support our interpretation of why, in the case of $N=3$, the adaptive players do not evolve toward the cooperative strategy $x^{(\mathrm{C})}$ but instead adopt the non-cooperative strategy $x^{(\mathrm{N})}$. To isolate the effect, we remove the autocratic player from the game and simulate a setting with only two adaptive agents. Payoff matrix is given in Table SI1, which has the same structure as the prisoner’s dilemma.
{adjustwidth}
-2.25in0in     $x^{(\mathrm{C})}$ $x^{(\mathrm{N})}$ $x^{(\mathrm{C})}$     $\frac{1}{8}\frac{(a-c)^{2}}{b},\frac{1}{8}\frac{(a-c)^{2}}{b}$ $\frac{5}{48}\frac{(a-c)^{2}}{b},\frac{5}{36}\frac{(a-c)^{2}}{b}$ $x^{(\mathrm{N})}$     $\frac{5}{36}\frac{(a-c)^{2}}{b},\frac{5}{48}\frac{(a-c)^{2}}{b}$ $\frac{1}{9}\frac{(a-c)^{2}}{b},\frac{1}{9}\frac{(a-c)^{2}}{b}$
Table SI1: Payoffs for only two agents. Both players take either $x^{(\mathrm{C})}$ or $x^{(\mathrm{N})}$.
Figures SI1 and SI2 show the results. In this case, we observe that both players eventually evolve to choose $x^{(\mathrm{N})}$ with probability one ($\sum q_{x^{(\mathrm{N})},x^{(\mathrm{N})}}/2=0$). Actually, in the $N=3$ setting, when the autocratic strategy takes $\overline{x}$, each adaptive player individually receives a higher payoff by choosing $x^{(\mathrm{N})}$ rather than $x^{(\mathrm{C})}$ (see Table 2).
As a consequence, the adaptive players quickly evolve strategies to respond to each other by selecting $x^{(\mathrm{N})}$, rather than adjusting their strategy to respond optimally to the autocratic player. This supports the finding that, unlike in the $N=2$ case, the autocratic strategy loses its power to promote cooperation when the number of adaptive players increases.
Fig SI1: Strategy adaptation of two adaptive memory-one players without an autocratic strategy.
Fig SI2: Time-averaged payoffs of two adaptive memory-one players without an autocratic strategy.
SI3 Averagely unbeatable zero-determinant alliance
In this appendix, we prove the existence of an averagely unbeatable zero-determinant (ZD) alliance. We assume that players $\mathcal{N}^{\prime}\subset\mathcal{N}:=\{1,\cdots,N\}$ with $\left|\mathcal{N}^{\prime}\right|=n$ $(1\leq n\leq N-1)$ adopt a ZD alliance. We write $\mathcal{N}^{\prime\prime}:=\mathcal{N}-\mathcal{N}^{\prime}$, $x_{\mathcal{N}^{\prime}}:=\sum_{k\in\mathcal{N}^{\prime}}x_{k}$, and $x_{\mathcal{N}^{\prime\prime}}:=\sum_{k\in\mathcal{N}^{\prime\prime}}x_{k}$. Similarly as Eq. (18), $x_{\mathcal{N}^{\prime}}$ and $x_{\mathcal{N}^{\prime\prime}}$ are effectively regarded as actions of single players, respectively. It should be noted that $x_{\mathcal{N}^{\prime}}\in[0,nx_{\mathrm{max}}]$ and $x_{\mathcal{N}^{\prime\prime}}\in[0,(N-n)x_{\mathrm{max}}]$.
We consider an averagely unbeatable ZD alliance, which unilaterally enforces
$\frac{1}{n}\sum_{k\in\mathcal{N}^{\prime}}\left\langle s_{k}\right\rangle^{*}$ $=\frac{1}{N-n}\sum_{k\in\mathcal{N}^{\prime\prime}}\left\langle s_{k}\right\rangle^{*}.$ (SI12)
For such case, we need to set $B$ as
$B\left(\bm{x}^{\prime}\right)$ $:=\frac{1}{n}\sum_{k\in\mathcal{N}^{\prime}}s_{k}\left(\bm{x}^{\prime}\right)-\frac{1}{N-n}\sum_{k\in\mathcal{N}^{\prime\prime}}s_{k}\left(\bm{x}^{\prime}\right).$ (SI13)
According to the Propositions, two actions satisfying
$B\left(\underline{x}_{\mathcal{N}^{\prime}},x_{\mathcal{N}^{\prime\prime}}\right)$ $\leq 0\quad(\forall x_{\mathcal{N}^{\prime\prime}})$
$B\left(\overline{x}_{\mathcal{N}^{\prime}},x_{\mathcal{N}^{\prime\prime}}\right)$ $\geq 0\quad(\forall x_{\mathcal{N}^{\prime\prime}})$ (SI14)
are necessary for the existence of a two-point averagely unbeatable ZD alliance. We rewrite $B$ as
$B\left(\bm{x}\right)$
$=\left[\left(a-c-b\sum_{l=1}^{N}x_{l}\right)\theta\left(a-b\sum_{l=1}^{N}x_{l}\right)-c\theta\left(-\left(a-b\sum_{l=1}^{N}x_{l}\right)\right)\right]\left[\frac{1}{n}x_{\mathcal{N}^{\prime}}-\frac{1}{N-n}x_{\mathcal{N}^{\prime\prime}}\right].$ (SI15)
First, we obtain
$B\left(\frac{n}{N}\frac{a-c}{b},x_{\mathcal{N}^{\prime\prime}}\right)$
$=\frac{b}{N-n}\left(\frac{N-n}{N}\frac{a-c}{b}-x_{\mathcal{N}^{\prime\prime}}\right)^{2}\theta\left(c+\frac{N-n}{N}(a-c)-bx_{\mathcal{N}^{\prime\prime}}\right)$
$\quad-\frac{c}{N-n}\left(\frac{N-n}{N}\frac{a-c}{b}-x_{\mathcal{N}^{\prime\prime}}\right)\theta\left(-\left(c+\frac{N-n}{N}(a-c)-bx_{\mathcal{N}^{\prime\prime}}\right)\right).$ (SI16)
The first term on the right-hand side is non-negative. When $c+(N-n)(a-c)/N-bx_{\mathcal{N}^{\prime\prime}}<0$, the second term on the right-hand side satisfies
$-\frac{c}{N-n}\left(\frac{N-n}{N}\frac{a-c}{b}-x_{\mathcal{N}^{\prime\prime}}\right)$ $>\frac{c}{N-n}\frac{c}{b}>0.$ (SI17)
Therefore, we obtain
$B\left(\frac{n}{N}\frac{a-c}{b},x_{\mathcal{N}^{\prime\prime}}\right)\geq 0.$ (SI18)
Second, we calculate
$B\left(nx_{\mathrm{max}},x_{\mathcal{N}^{\prime\prime}}\right)$ $=-c\left[x_{\mathrm{max}}-\frac{1}{N-n}x_{\mathcal{N}^{\prime\prime}}\right]\leq 0.$ (SI19)
Therefore, we can construct a two-point averagely unbeatable ZD alliance by using $\underline{x}_{\mathcal{N}^{\prime}}=nx_{\mathrm{max}}$ and $\overline{x}_{\mathcal{N}^{\prime}}=n(a-c)/(Nb)$.
Acknowledgments
M.U. was supported by JSPS KAKENHI Grant Number JP20K19884 and Toyota Riken Scholar Program. G.I. was supported by JSPS KAKENHI Grant Number JP23K22982.